Two cells of emf 1v, 2v and internal resistance 2volt per meter respectively connected serious and parallel.  What should be the external resistance be the same
 

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please repost the question with understandable english so that it could be solved
Or please upload the source or image of question

Regards
Khimraj
4 years ago
In series effective emf =2+1=3V
effective internal resistance =2+1=3ohm 
total resistance =3+R ohm 
where R=external resistance
current through the circuit=current through the external resistance =3/3+R ampere 
in parallel effective emf =2V
effective resistance =2/3ohm 
total resistance =2/3+R. ohm
current through the external resistance
=2/(2/3+R) Ampere
in two cases current is same 
then solving two equation we get R=4ohm
as in series connection resistance is more so heat will generate more in series connection 
hope this will help you